# Deep Positive-Unlabeled Anomaly Detection for Contaminated Unlabeled Data

### v1

@@ Abstract
Semi-supervised anomaly detection, which aims to improve the performance of the anomaly detector by using a small amount of anomaly data in addition to unlabeled data, has attracted attention. Existing semi-supervised approaches assume that unlabeled data are mostly normal. They train the anomaly detector to minimize the anomaly scores for the unlabeled data, and to maximize those for the anomaly data. However, in practice, the unlabeled data are often contaminated with anomalies. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the detection performance. To solve this problem, we propose the positive-unlabeled autoencoder, which is based on positive-unlabeled learning and the anomaly detector such as the autoencoder. With our approach, we can approximate the anomaly scores for normal data using the unlabeled and anomaly data. Therefore, without the labeled normal data, we can train the anomaly detector to minimize the anomaly scores for normal data, and to maximize those for the anomaly data. In addition, our approach is applicable to various anomaly detectors such as the DeepSVDD. Experiments on various datasets show that our approach achieves better detection performance than existing approaches.
@@ 1 Introduction
Anomaly detection, which aims to identify unusual data points, is an important task in machine learning [48]. It has been performed in various fields [24, 11, 28, 1, 26, 3, 30, 4, 42], such as cyber-security, fault and damage detection, novelty detection, fraud detection, and medical diagnosis.
In general, anomaly detection is performed by unsupervised learning because it does not require expensive and time-consuming labeling. Unsupervised anomaly detection assumes that unlabeled data are mostly normal, and tries to detect anomalies by using the difference from them. As an example, we focus on the autoencoder (AE) [16], which has been successfully applied to anomaly detection [50]. The AE learns the low-dimensional representation of data points through data reconstruction. The AE consists of two neural networks: the encoder and the decoder. The encoder maps a data point to the representation, and the decoder reconstructs the original data point from the representation. The AE is trained to minimize the reconstruction error, which is the difference between the data point and its reconstructed data point. After training, the AE is expected to successfully reconstruct normal data and fail to reconstruct anomaly data because the training dataset is assumed to contain only normal data and no anomaly data. Hence, the reconstruction error can be used for the anomaly score. Although unsupervised approaches are often used in anomaly detection, they have limitations in detection performance because they cannot use information about anomalies.
@@
To improve the anomaly detection performance, semi-supervised anomaly detection uses a small amount of anomaly data in addition to unlabeled data. Existing approaches train the anomaly detector to minimize the anomaly scores for the unlabeled data, and to maximize those for the anomaly data [15, 47, 59]. For example, the autoencoding binary classifier (ABC) [59] trains the AE to minimize the reconstruction errors for the unlabeled data, and to maximize those for the anomaly data. However, in practice, the unlabeled data are often contaminated with anomalies. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the anomaly detection performance. This frequently occurs because it is difficult to label all anomalies.
@@
Our purpose is to propose a semi-supervised learning approach that can improve the anomaly detection performance even if the unlabeled data are contaminated with anomalies. To handle such unlabeled data, we propose the positive-unlabeled autoencoder (PUAE), which is based on positive-unlabeled (PU) learning [9, 8, 22] and the anomaly detector such as the AE. PU learning assumes that an unlabeled data distribution is a mixture of normal and anomaly data distributions111 Note that the anomaly data in the training dataset follow the anomaly data distribution, but new types of anomalies, not included in the training dataset, do NOT follow this distribution. In general, no distribution can fully represent all possible anomalies. . Accordingly, the normal data distribution is approximated by the unlabeled and anomaly data distributions. With this assumption, we approximate the reconstruction errors for normal data with the unlabeled and anomaly data. Therefore, without the labeled normal data, we can train the AE to minimize the reconstruction errors for normal data, and to maximize those for the anomaly data.
@@
Figure 1 shows the difference between the PU learning [22], the denoising AE (DAE) [55], the ABC, and the proposed method on the toy dataset. The DAE is a variant of the AE, which tries to reconstruct original data points from noisy input data points. This has also been successfully applied to anomaly detection [50]. Here, the ABC and the proposed method are based on the DAE. The toy dataset consists of the unlabeled and anomaly data, where the unlabeled data include both normal and anomaly data points. We first focus on the PU learning, which aims to train the binary classifier from the unlabeled and anomaly data. This can detect seen anomalies that are included in the training dataset. However, since its decision boundary is between normal data points and seen anomalies, it cannot detect unseen anomalies, which are new types of anomalies unseen during training, such as novel anomalies and zero-day attacks [56, 36, 7]. We next focus on the DAE and the ABC. They can detect unseen anomalies to some extent since they try to detect anomalies by using the difference from normal data. The DAE cannot detect seen anomalies because it cannot use information about anomalies. The ABC can detect seen anomalies to some extent since it can use the anomaly data. However, the contaminated dataset weakens the effect of maximizing the anomaly scores for anomalies in the ABC. Finally, we focus on the proposed method. The proposed method can detect seen anomalies according to the effectiveness of PU learning, and can detect unseen anomalies to some extent according to the effectiveness of the DAE.
Furthermore, our approach is applicable to various anomaly detectors. When choosing the detector, our approach requires that its loss function be non-negative and differentiable. A lot of approaches satisfy this requirement, such as the deep support vector data description (DeepSVDD) [46]. Our approaches achieve better detection performance than existing approaches on various datasets.
@@
  • We propose the positive-unlabeled autoencoder (PUAE), the semi-supervised anomaly detection approach that can handle contaminated unlabeled data.
  • Our approach is applicable to various anomaly detectors such as the DeepSVDD.
  • Experiments on various datasets show that our approach achieves better anomaly detection performance than existing approaches.
@@ 2 Preliminaries
In this section, we first explain our problem setup. Next, we review the autoencoder (AE) [16] and the autoencoding binary classifier (ABC) [59]. They are typical unsupervised and semi-supervised anomaly detection approaches, and our approach is based on them.
@@ 2.1 Problem Setup
Given unlabeled dataset $\mathcal{U}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ and anomaly dataset $\mathcal{A}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{M}\}$ for training. $\mathcal{U}$ contains not only normal data points but also seen anomalies included in $\mathcal{A}$. The test dataset contains normal data points, seen anomalies, and unseen anomalies not included in $\mathcal{A}$. Our goal is to obtain a high performance anomaly detector using $\mathcal{U}$ and $\mathcal{A}$.
@@ 2.2 Autoencoder
For unsupervised anomaly detectors, we train them only using the dataset $\mathcal{U}$. As an example, we focus on the AE, which has been successfully applied to anomaly detection [50]. The AE is presented for representation learning, which learns the representation of data points through data reconstruction. Let $\mathbf{x}$ be a data point and $\mathbf{z}$ be its low-dimensional latent representation. The AE consists of two neural networks: encoder $E_{\theta}(\mathbf{x})$ and decoder $D_{\theta}(\mathbf{z})$, where $\theta$ is the parameter of these neural networks. $E_{\theta}(\mathbf{x})$ maps a data point $\mathbf{x}$ into a low-dimensional latent representation $\mathbf{z}$, and $D_{\theta}(\mathbf{z})$ reconstructs the original data point $\mathbf{x}$ from the latent representation $\mathbf{z}$. The reconstruction error for each data point $\mathbf{x}$ in the AE is defined as follows:
@@
Although unsupervised approaches are often used in anomaly detection, they have limitations in detection performance because they cannot use information about anomalies.
@@ 2.3 Autoencoding Binary Classifier
Semi-supervised anomaly detection aims to improve the anomaly detection performance by using the datasets $\mathcal{U}$ and $\mathcal{A}$. A number of studies have been presented such as the ABC [59], the deep semi-supervised anomaly detection (DeepSAD) [47], and the outlier exposure [15]. Here, we focus on the ABC, which is based on the AE.
Let $y=0$ be normal and $y=1$ be anomaly. The ABC models the conditional probability of the normality or the abnormality $y$ given the data point $\mathbf{x}$ by using the reconstruction error $\ell(\mathbf{x};\theta)$ as follows:
@@
A small reconstruction error results in a higher $p_{\theta}(y=0|\mathbf{x})$, while a large reconstruction error results in a higher $p_{\theta}(y=1|\mathbf{x})$.
@@
Like the AE, the ABC assumes all unlabeled data points to be normal. We train the ABC by minimizing the following objective function:
@@
This minimizes the reconstruction errors for the unlabeled data and maximizes those for the anomaly data. Hence, after training, the AE becomes to reconstruct the unlabeled data assumed to be normal, and fail to reconstruct anomalies. Other semi-supervised anomaly detection approaches such as the DeepSAD [47] and the outlier exposure [15] also minimize the anomaly scores for the unlabeled data and maximize those for the anomaly data.
However, the unlabeled dataset $\mathcal{U}$ is often contaminated with anomalies in practice. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the anomaly detection performance. This frequently occurs because it is difficult to label all anomalies.
@@ 3 Proposed Method
We aim to improve the detection performance even if the unlabeled dataset $\mathcal{U}$ contains anomalies. To handle such unlabeled dataset, we propose the positive-unlabeled autoencoder (PUAE), which is based on positive-unlabeled (PU) learning [9, 8, 22] and the anomaly detector such as the AE. Hereinafter, we also refer anomalies as positive (+) samples, and normal data points as negative (-) samples.
@@ 3.1 Positive-Unlabeled Autoencoder
Let $p_{\mathcal{N}}$ be the normal data distribution, $p_{\mathcal{A}}$ be the seen anomaly distribution, and $p_{\mathcal{U}}$ be the unlabeled data distribution. We assume that the datasets $\mathcal{U}$ and $\mathcal{A}$ are drawn from $p_{\mathcal{U}}$ and $p_{\mathcal{A}}$, respectively. We also assume that $p_{\mathcal{U}}$ can be rewritten as follows:
@@
where $\alpha\in[0,1]$ represents the probability of anomaly occurrence in the unlabeled data. Hence, $p_{\mathcal{N}}$ can be rewritten as follows:
@@
Although $\alpha$ is the hyperparameter and is assumed to be known throughout this paper, it can be estimated from the datasets $\mathcal{U}$ and $\mathcal{A}$ in conventional PU learning approaches [29, 44, 17, 6].
@@
Hence, using the seen anomaly distribution $p_{\mathcal{A}}$ and the unlabeled data distribution $p_{\mathcal{U}}$, $\mathcal{L}_{\mathrm{PN}}(\theta)$ can be rewritten as follows:
@@
With the datasets $\mathcal{U}$ and $\mathcal{A}$, we can approximate $\mathcal{L}_{\mathrm{PN}}(\theta)$ by the Monte Carlo approximation as follows:
@@
The left-hand side in Eq. (12) is always greater than or equal to zero, but the right-hand side can be negative. In experiments, it often converges towards negative infinity, resulting in the meaningless solution. To avoid this, according to [22], our training objective function to be minimized ensures that $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)$ is not negative as follows:
$\mathcal{L}_{\mathrm{PUAE}}(\theta)=\alpha\mathcal{L}_{\mathcal{A}}^{+}(\theta)+\max\left\{0,\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)\right\}.$ (13)
@@
1:Unlabeled and anomaly datasets $(\mathcal{U},\mathcal{A})$, mini-batch size $K$, hyperparameter $\alpha\in[0,1]$
2:Model parameter $\theta$
3:while not converged do
4:     Sample mini-batch $\mathcal{B}$ from datasets $(\mathcal{U},\mathcal{A})$
5:     Compute $\mathcal{L}_{\mathcal{A}}^{+}(\theta)$, $\mathcal{L}_{\mathcal{U}}^{-}(\theta)$, and $\mathcal{L}_{\mathcal{A}}^{-}(\theta)$ in Eq. (11) with $\mathcal{B}$
6:     if $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)\geq 0$ then
7:         Set the gradient $\nabla_{\theta}(\alpha\mathcal{L}_{\mathcal{A}}^{+}(\theta)+\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta))$
8:     else
9:         Set the gradient $\nabla_{\theta}-(\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta))$
10:     end if
11:     Update $\mathbf{\theta}$ with the gradient
12:end while
We can optimize this training objective function by using the stochastic gradient descent (SGD) such as Adam [19]. In practice, according to [22], when $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)<0$, we maximize it to be greater than or equal to zero. We call our proposed method the positive-unlabeled autoencoder (PUAE). Algorithm 1 shows the pseudo code of the PUAE, where $K$ is the mini-batch size for the SGD. Note that our approach can be easily extended to positive-negative-unlabeled (PNU) learning [49], where we can also use a small amount of normal data.

### v2

@@ Abstract
Semi-supervised anomaly detection, which aims to improve the anomaly detection performance by using a small amount of labeled anomaly data in addition to unlabeled data, has attracted attention. Existing semi-supervised approaches assume that most unlabeled data are normal, and train anomaly detectors by minimizing the anomaly scores for the unlabeled data while maximizing those for the labeled anomaly data. However, in practice, the unlabeled data are often contaminated with anomalies. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the detection performance. To solve this problem, we propose the deep positive-unlabeled anomaly detection framework, which integrates positive-unlabeled learning with deep anomaly detection models such as autoencoders and deep support vector data descriptions. Our approach enables the approximation of anomaly scores for normal data using the unlabeled data and the labeled anomaly data. Therefore, without labeled normal data, our approach can train anomaly detectors by minimizing the anomaly scores for normal data while maximizing those for the labeled anomaly data. Experiments on various datasets show that our approach achieves better detection performance than existing approaches.
Keywords Semi-Supervised Anomaly Detection, Positive-Unlabeled Learning
@@ 1 Introduction
Anomaly detection, which aims to identify unusual data points, is an important task in machine learning (Ruff et al., 2021). It has been performed in various fields such as cyber-security (Kwon et al., 2019), novelty detection (Marchi et al., 2015), medical diagnosis (Litjens et al., 2017), infrastructure monitoring (Borghesi et al., 2019), and natural sciences (Min et al., 2017; Cerri et al., 2019; Pracht et al., 2020).
In general, anomaly detection is performed by unsupervised learning because it does not require expensive and time-consuming labeling. Unsupervised approaches assume that most unlabeled data are normal, and try to detect anomalies by using an anomaly score, which represents the difference from normal data (Hinton and Salakhutdinov, 2006; Ruff et al., 2018). Although these approaches are easy to handle, their detection performance is limited because they cannot use information about anomalies.
@@
To improve the detection performance, semi-supervised anomaly detection uses a small amount of labeled anomaly data in addition to unlabeled data. Existing semi-supervised approaches train anomaly detectors to minimize the anomaly scores for the unlabeled data, and to maximize those for the labeled anomaly data (Hendrycks et al., 2018; Ruff et al., 2019; Yamanaka et al., 2019). However, in practice, the unlabeled data are often contaminated with anomalies. This weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the anomaly detection performance. This frequently occurs because it is difficult to label all anomalies.
@@
Our purpose is to propose a semi-supervised approach that can improve the anomaly detection performance even if the unlabeled data are contaminated with anomalies. To handle such unlabeled data, we propose the deep positive-unlabeled anomaly detection framework, which integrates positive-unlabeled (PU) learning (Du Plessis et al., 2014, 2015; Kiryo et al., 2017) with deep anomaly detectors such as the autoencoder (AE) (Hinton and Salakhutdinov, 2006) and the deep support vector data description (DeepSVDD) (Ruff et al., 2018). PU learning assumes that an unlabeled data distribution is a mixture of normal and anomaly data distributions111 Note that the anomaly data in the training dataset follow the anomaly data distribution, but new types of anomalies, unseen during training, may NOT follow this distribution. In general, no distribution can fully represent all possible anomalies. . Accordingly, the normal data distribution is approximated by using the unlabeled and anomaly data distributions. With this assumption, we approximate the anomaly scores for normal data using the unlabeled data and the labeled anomaly data. Therefore, without labeled normal data, we can train anomaly detectors to minimize the anomaly scores for normal data, and to maximize those for the labeled anomaly data.
@@
Figure 1 compares the PU learning (Kiryo et al., 2017), the unsupervised detector, the semi-supervised detector, and our approach on the toy dataset. We used the denoising AE (DAE) (Vincent et al., 2008) for the unsupervised detector, and the autoencoding binary classifier (ABC) (Yamanaka et al., 2019) for the semi-supervised detector. Our approach is based on the DAE. The toy dataset consists of unlabeled and anomaly data, where the unlabeled data include both normal and anomaly data points. We first focus on the PU learning, which aims to train the binary classifier from the unlabeled data and the labeled anomaly data. This can detect seen anomalies, which are similar to anomalies included in the training dataset. However, since its decision boundary is between normal data points and seen anomalies, it cannot detect unseen anomalies, which are new types of anomalies unseen during training, such as novel anomalies and zero-day attacks (Wang et al., 2013; Pang et al., 2021; Ding et al., 2022). We next focus on unsupervised and semi-supervised detectors. They can detect unseen anomalies to some extent since they try to detect anomalies by using the difference from normal data. The unsupervised detector cannot detect seen anomalies since it cannot use information about anomalies. The semi-supervised detector can detect seen anomalies to some extent since it can use the labeled anomaly data. However, the contaminated dataset weakens the effect of maximizing the anomaly scores for anomalies in the semi-supervised detector. Finally, we focus on our approach. our approach can detect seen anomalies according to the effectiveness of PU learning, and can detect unseen anomalies to some extent according to the effectiveness of the deep anomaly detector.
Our framework is applicable to various anomaly detector. When selecting a detector, we require that its loss function be non-negative and differentiable. For example, the AE, the DeepSVDD, and recent self-supervised detectors such as (Hendrycks et al., 2019; Qiu et al., 2021; Shenkar and Wolf, 2021) satisfy this. In this paper, we apply our framework to the AE and the DeepSVDD. We refer to the former as the positive-unlabeled autoencoder (PUAE), and the latter as the positive-unlabeled support vector data description (PUSVDD).
Compared to existing semi-supervised approaches designed to handle contaminated unlabeled data (Zhang et al., 2018; Ju et al., 2020; Zhang et al., 2021; Pang et al., 2023; Li et al., 2023; Perini et al., 2023), the proposed method is theoretically justified from the perspective of unbiased PU learning (Du Plessis et al., 2014, 2015; Kiryo et al., 2017). In addition, the proposed method achieved equal to or better performance than the current state-of-the-art approach (Li et al., 2023) across eight datasets.
@@
  • To handle contaminated unlabeled data, we propose the deep positive-unlabeled anomaly detection framework, which integrates unbiased PU learning with the deep anomaly detectors such as the AE and the DeepSVDD.
  • Experiments on various datasets demonstrate that our approach outperforms existing approaches in anomaly detection performance.
@@ 2 Preliminaries
In this section, we first explain our problem setup. Next, we review the AE (Hinton and Salakhutdinov, 2006) and the ABC (Yamanaka et al., 2019). They are typical unsupervised and semi-supervised anomaly detection approaches, and our framework can be applied to them.
@@ 2.1 Problem Setup
Given unlabeled dataset $\mathcal{U}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$ and anomaly dataset $\mathcal{A}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{M}\}$ for training. $\mathcal{U}$ contains not only normal data points but also seen anomalies that are similar to $\mathcal{A}$. The test dataset contains normal data points, seen anomalies, and unseen anomalies that are new types of anomalies unseen during training. Our goal is to obtain a high performance anomaly detector using $\mathcal{U}$ and $\mathcal{A}$.
@@ 2.2 Autoencoder
For unsupervised anomaly detectors, we train them only using the unlabeled dataset $\mathcal{U}$. As an example, we focus on the AE, which has been successfully applied to anomaly detection (Sakurada and Yairi, 2014). The AE is presented for representation learning, which learns the representation of data points through data reconstruction. Let $\mathbf{x}$ be a data point and $\mathbf{z}$ be its low-dimensional latent representation. The AE consists of two neural networks: encoder $E_{\theta}(\mathbf{x})$ and decoder $D_{\theta}(\mathbf{z})$, where $\theta$ is the parameter of these neural networks. $E_{\theta}(\mathbf{x})$ maps a data point $\mathbf{x}$ into a low-dimensional latent representation $\mathbf{z}$, and $D_{\theta}(\mathbf{z})$ reconstructs the original data point $\mathbf{x}$ from the latent representation $\mathbf{z}$. The reconstruction error for each data point $\mathbf{x}$ in the AE is defined as follows:
@@
Although unsupervised approaches are widely used in anomaly detection, their detection performance is limited because they cannot use information about anomalies.
@@ 2.3 Autoencoding Binary Classifier
Semi-supervised anomaly detection aims to improve the anomaly detection performance using the unlabeled dataset $\mathcal{U}$ and the anomaly dataset $\mathcal{A}$. A number of studies have been presented such as the ABC (Yamanaka et al., 2019), the deep semi-supervised anomaly detection (DeepSAD) (Ruff et al., 2019), and the outlier exposure (Hendrycks et al., 2018). Here, we focus on the ABC, which is based on the AE.
Let $y=0$ be normal and $y=1$ be anomaly. The ABC models the conditional probability of $y$ given $\mathbf{x}$ by using the reconstruction error $\ell(\mathbf{x};\theta)$ as follows:
@@
A small reconstruction error results in a higher probability of normality $p_{\theta}(y=0|\mathbf{x})$, while a large reconstruction error results in a higher probability of abnormality $p_{\theta}(y=1|\mathbf{x})$.
@@
Like the AE, the ABC assumes all unlabeled data points to be normal. The ABC is trained by minimizing the following objective function:
@@
This minimizes the reconstruction errors for the unlabeled data and maximizes those for the anomaly data. Hence, after training, the AE becomes to reconstruct the unlabeled data assumed to be normal, and fail to reconstruct anomalies. Other semi-supervised anomaly detection approaches such as the DeepSAD (Ruff et al., 2019) and the outlier exposure (Hendrycks et al., 2018) also minimize the anomaly scores for the unlabeled data and maximize those for the anomaly data.
However, the unlabeled dataset $\mathcal{U}$ is often contaminated with anomalies in practice. This contamination weakens the effect of maximizing the anomaly scores for anomalies, and prevents us from improving the detection performance. This frequently occurs because it is difficult to label all anomalies in the unlabeled dataset.
@@ 3 Proposed Method
We aim to improve the detection performance even if the unlabeled dataset $\mathcal{U}$ contains anomalies. To handle such unlabeled dataset, we propose the deep positive-unlabeled anomaly detection framework, which integrates PU learning (Du Plessis et al., 2014, 2015; Kiryo et al., 2017) with deep anomaly detection models such as the AE, the DeepSVDD, and recent self-supervised detectors such as (Hendrycks et al., 2019; Qiu et al., 2021; Shenkar and Wolf, 2021).
In this section, we apply our framework to the AE and the DeepSVDD. We refer to the former as the positive-unlabeled autoencoder (PUAE), and the latter as the positive-unlabeled support vector data description (PUSVDD).
Hereinafter, we also refer anomalies as positive (+) samples, and normal data points as negative (-) samples.
@@ 3.1 Positive-Unlabeled Autoencoder
At first, we explain the PUAE. Let $p_{\mathcal{N}}$ be the normal data distribution, $p_{\mathcal{A}}$ be the seen anomaly distribution, and $p_{\mathcal{U}}$ be the unlabeled data distribution. We assume that the datasets $\mathcal{U}$ and $\mathcal{A}$ are drawn from $p_{\mathcal{U}}$ and $p_{\mathcal{A}}$, respectively. We also assume that $p_{\mathcal{U}}$ can be rewritten as follows:
@@
where $\alpha\in[0,1]$ is the probability of anomaly occurrence in the unlabeled data. Hence, $p_{\mathcal{N}}$ can be rewritten as follows:
@@
Although $\alpha$ is the hyperparameter and is assumed to be known throughout this paper, it can be estimated from the datasets $\mathcal{U}$ and $\mathcal{A}$ in conventional PU learning approaches (Menon et al., 2015; Ramaswamy et al., 2016; Jain et al., 2016; Christoffel et al., 2016).
@@
Hence, by using the seen anomaly distribution $p_{\mathcal{A}}$ and the unlabeled data distribution $p_{\mathcal{U}}$, $\mathcal{L}_{\mathrm{PN}}(\theta)$ can be rewritten as follows:
@@
With the datasets $\mathcal{U}$ and $\mathcal{A}$, we can approximate $\mathcal{L}_{\mathrm{PN}}(\theta)$ by the empirical distribution as follows:
@@
The left-hand side in Eq. (12) is always greater than or equal to zero, but the right-hand side can be negative. In experiments, it often converges towards negative infinity, resulting in the meaningless solution. To avoid this, based on (Hammoudeh and Lowd, 2020), our training objective function to be minimized ensures that $\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)$ is not negative as follows:
$\mathcal{L}_{\mathrm{Proposed}}(\theta)=\alpha\mathcal{L}_{\mathcal{A}}^{+}(\theta)+\left|\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)\right|.$ (13)
We can optimize this training objective function by using the stochastic gradient descent (SGD) such as Adam (Kingma and Ba, 2015). We refer to this approach as the PUAE.
Algorithm 1 shows the pseudo code of the PUAE, where $K$ is the mini-batch size for the SGD. Note that our approach can be easily extended to positive-negative-unlabeled (PNU) learning (Sakai et al., 2017), where we can also use a small amount of labeled normal data.
@@
Input: Unlabeled and anomaly datasets $(\mathcal{U},\mathcal{A})$, mini-batch size $K$, hyperparameter $\alpha\in[0,1]$
Output: Model parameter $\theta$
Procedure:
1:  while not converged do
2:     Sample mini-batch $\mathcal{B}$ from datasets $(\mathcal{U},\mathcal{A})$
3:     Compute $\mathcal{L}_{\mathcal{A}}^{+}(\theta)$, $\mathcal{L}_{\mathcal{U}}^{-}(\theta)$, and $\mathcal{L}_{\mathcal{A}}^{-}(\theta)$ in Eq. (11) with $\mathcal{B}$
4:     Set the gradient $\nabla_{\theta}(\alpha\mathcal{L}_{\mathcal{A}}^{+}(\theta)+\left|\mathcal{L}_{\mathcal{U}}^{-}(\theta)-\alpha\mathcal{L}_{\mathcal{A}}^{-}(\theta)\right|)$
5:     Update $\mathbf{\theta}$ with the gradient
6:  end while
3.2 Positive-Unlabeled Support Vector Data Description
We next apply our framework to the DeepSVDD. The DeepSVDD aims to pull the representation of the normal data towards the pre-defined center, and push those of the anomaly data away from the center. Let $f_{\theta}(\mathbf{x})$ be the feature extractor, like the encoder in the AE. The loss function for each data point of the DeepSVDD is defined as follows:
~$\tilde{\ell}(\mathbf{x};\theta)=\left\|f_{\theta}(\mathbf{x})-\mathbf{c}\right\|^{2},$ (14)
where $\mathbf{c}\neq\mathbf{0}$ is the pre-defined center vector.
The DeepSAD (Ruff et al., 2019) is a semi-supervised extension of the DeepSVDD. The DeepSAD trains the DeepSVDD model to minimize Eq. (14) for the unlabeled data, and to maximize it for the anomaly data. The loss function for each data point of the DeepSAD is defined as follows:
~$\tilde{\ell}_{\mathrm{SAD}}(\mathbf{x},y;\theta)=(1-y)\tilde{\ell}(\mathbf{x};\theta)+\frac{y}{\tilde{\ell}(\mathbf{x};\theta)}.$ (15)
We can apply our framework to the DeepSVDD by replacing Eq. (4) in the PUAE with Eq. (15), while keeping all other components identical to the PUAE. We refer to this approach as the PUSVDD.